OpenStudy (anonymous):

Express in terms of i. 4√-225

OpenStudy (saifoo.khan):

60 i

OpenStudy (anonymous):

Can you show your work, please?

OpenStudy (zarkon):

$$\sqrt{225}=15$$ $$4\cdot 15=60$$ since $$\sqrt{-1}=i$$ we get $$60i$$